OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 4(x - 6) = -2(x - 3)

OpenStudy (anonymous):

expand the brackets 4x-24=-2x + 6 (-ve time -ve= +ve) 4x +2x= 6 +24 6x = 30 x= 30/6 x= 5

OpenStudy (anonymous):

times*